### v2

@@ Abstract
Robot-assisted minimally invasive surgery benefits from enhancing dynamic scene reconstruction, as it improves surgical outcomes. While Neural Radiance Fields (NeRF) have been effective in scene reconstruction, their slow inference speeds and lengthy training durations limit their applicability. To overcome these limitations, 3D Gaussian Splatting (3D-GS) based methods have emerged as a recent trend, offering rapid inference capabilities and superior 3D quality. However, these methods still struggle with under-reconstruction in both static and dynamic scenes. In this paper, we propose HFGS, a novel approach for deformable endoscopic reconstruction that addresses these challenges from spatial and temporal frequency perspectives. Our approach incorporates deformation fields to better handle dynamic scenes and introduces Spatial High-Frequency Emphasis Reconstruction (SHF) to minimize discrepancies in spatial frequency spectra between the rendered image and its ground truth. Additionally, we introduce Temporal High-Frequency Emphasis Reconstruction (THF) to enhance dynamic awareness in neural rendering by leveraging flow priors, focusing optimization on motion-intensive parts. Extensive experiments on two widely used benchmarks demonstrate that HFGS achieves superior rendering quality. Our code will be available.
@@ 1 Introduction
Endoscopic procedures are foundational to minimally invasive surgery, significantly reducing trauma and hastening patient recovery [Gao et al.(2024)Gao, Yang, Xiao, Zhu, Zhang, Hou, Liu, Meng, Sun, Zuo, et al., Psychogyios et al.(2023)Psychogyios, Colleoni, Van Amsterdam, Li, Huang, Li, Jia, Zou, Wang, Liu, et al.]. In Robotic-Assisted Minimally Invasive Surgery (RAMIS), the reconstruction of a 3D model of the surgical scene from stereo endoscopes is critical for surgical precision and efficiency. This technology enables surgeons to visualize observed tissues in 3D, which enhances their spatial awareness and navigation capabilities [Mahmoud et al.(2017)Mahmoud, Cirauqui, Hostettler, Doignon, Soler, Marescaux, and Montiel]. Despite the many benefits of endoscopic reconstruction, several challenges remain, including limited field-of-view, obstructions, and dynamic tissue deformation [Scott et al.(2008)Scott, Cendan, Pugh, Minter, Dunnington, and Kozar, Shin et al.(2019)Shin, Ferguson, Pedram, Ma, Dutson, and Rosen, Zha et al.(2023)Zha, Cheng, Li, Harandi, and Ge, Yang et al.(2023b)Yang, Wang, Wang, Yang, and Shen, Wang et al.(2022)Wang, Long, Fan, and Dou]. Previous studies [Zha et al.(2023)Zha, Cheng, Li, Harandi, and Ge, Yang et al.(2023b)Yang, Wang, Wang, Yang, and Shen, Wang et al.(2022)Wang, Long, Fan, and Dou, Huang et al.(2024)Huang, Cui, Bai, Guo, Xu, and Ren, Liu et al.(2024)Liu, Li, Yang, and Yuan, Zhu et al.(2024)Zhu, Wang, Jin, Lin, and Yu] have successfully employed depth maps for endoscopic reconstruction; however, these methods still face two significant issues: the lack of sufficient details in generated models and inadequate rendering of non-rigid deformations.
@@
To address these issues, 3D Gaussian Splatting (3D-GS) [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] emerges as an effective alternative, providing rapid inference capability and enhanced quality of 3D representation. By optimizing anisotropic 3D Gaussians with a collection of scene images, 3D-GS effectively captures the spatial positioning, orientations, color properties, and alpha blending factors. 3D-GS reconstructs not only the geometry but also the visual texture of scenes with rapid rendering performances [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis]. Although 3D-GS is extended to represent dynamic scenes [Zhu et al.(2024)Zhu, Wang, Jin, Lin, and Yu, Liu et al.(2024)Liu, Li, Yang, and Yuan, Huang et al.(2024)Huang, Cui, Bai, Guo, Xu, and Ren, Katsumata et al.(2023)Katsumata, Vo, and Nakayama, Yang et al.(2024b)Yang, Yang, Pan, and Zhang, Luiten et al.(2024)Luiten, Kopanas, Leibe, and Ramanan], it often suffers from under-reconstruction [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] during the process of Gaussian densification [Zhang et al.(2024)Zhang, Zhan, Xu, Lu, and Xing], which affects both static and dynamic scenes. The under-reconstruction can be clearly observed with blur and artifacts in the rendered 2D images, the discrepancy of frequency spectrum of the render images and the corresponding ground truth, and the predicted optical flow results by [Xu et al.(2023)Xu, Zhang, Cai, Rezatofighi, Yu, Tao, and Geiger], as illustrated in Fig. 1.

### v3

@@ Abstract
Robot-assisted minimally invasive surgery benefits from enhancing dynamic scene reconstruction, as it improves surgical outcomes. While Neural Radiance Fields (NeRF) have been effective in scene reconstruction, their slow inference speeds and lengthy training durations limit their applicability. To overcome these limitations, 3D Gaussian Splatting (3D-GS) based methods have emerged as a recent trend, offering rapid inference capabilities and superior 3D quality. However, these methods still struggle with under-reconstruction in both static and dynamic scenes. In this paper, we propose HFGS, a novel approach for deformable endoscopic reconstruction that addresses these challenges from spatial and temporal frequency perspectives. Our approach incorporates deformation fields to better handle dynamic scenes and introduces Spatial High-Frequency Emphasis Reconstruction (SHF) to minimize discrepancies in spatial frequency spectra between the rendered image and its ground truth. Additionally, we introduce Temporal High-Frequency Emphasis Reconstruction (THF) to enhance dynamic awareness in neural rendering by leveraging flow priors, focusing optimization on motion-intensive parts. Extensive experiments on two widely used benchmarks demonstrate that HFGS achieves superior rendering quality. Source code is available at https://github.com/zhaohaoyu376/HFGS.
@@ 1 Introduction
Endoscopic procedures are foundational to minimally invasive surgery, significantly reducing trauma and hastening patient recovery [Gao et al.(2024)Gao, Yang, Xiao, Zhu, Zhang, Hou, Liu, Meng, Sun, Zuo, et al., Psychogyios et al.(2023)Psychogyios, Colleoni, Van Amsterdam, Li, Huang, Li, Jia, Zou, Wang, Liu, et al., Wang et al.(2023)Wang, Jin, and Zhu]. In Robotic-Assisted Minimally Invasive Surgery (RAMIS), the reconstruction of a 3D model of the surgical scene from stereo endoscopes is critical for surgical precision and efficiency. This technology enables surgeons to visualize observed tissues in 3D, which enhances their spatial awareness and navigation capabilities [Mahmoud et al.(2017)Mahmoud, Cirauqui, Hostettler, Doignon, Soler, Marescaux, and Montiel, Wang et al.(2024)Wang, Zhu, Yang, Guo, Zhang, Xu, and Jin]. Despite the many benefits of endoscopic reconstruction, several challenges remain, including limited field-of-view, obstructions, and dynamic tissue deformation [Scott et al.(2008)Scott, Cendan, Pugh, Minter, Dunnington, and Kozar, Zha et al.(2023)Zha, Cheng, Li, Harandi, and Ge, Yang et al.(2023b)Yang, Wang, Wang, Yang, and Shen, Wang et al.(2022)Wang, Long, Fan, and Dou]. Previous studies [Zha et al.(2023)Zha, Cheng, Li, Harandi, and Ge, Yang et al.(2023b)Yang, Wang, Wang, Yang, and Shen, Wang et al.(2022)Wang, Long, Fan, and Dou, Huang et al.(2024)Huang, Cui, Bai, Guo, Xu, and Ren, Liu et al.(2024)Liu, Li, Yang, and Yuan, Zhu et al.(2024)Zhu, Wang, Jin, Lin, and Yu] have successfully employed depth maps for endoscopic reconstruction; however, these methods still face two significant issues: the lack of sufficient details in generated models and inadequate rendering of non-rigid deformations.
@@
To address these issues, 3D Gaussian Splatting (3D-GS) [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] emerges as an effective alternative, providing rapid inference capability and enhanced quality of 3D representation. By optimizing anisotropic 3D Gaussians with a collection of scene images, 3D-GS effectively captures the spatial positioning, orientations, color properties, and alpha blending factors. 3D-GS reconstructs not only the geometry but also the visual texture of scenes with rapid rendering performances [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis]. Although 3D-GS is extended to represent dynamic scenes [Zhu et al.(2024)Zhu, Wang, Jin, Lin, and Yu, Liu et al.(2024)Liu, Li, Yang, and Yuan, Huang et al.(2024)Huang, Cui, Bai, Guo, Xu, and Ren, Yang et al.(2024b)Yang, Yang, Pan, and Zhang, Luiten et al.(2024)Luiten, Kopanas, Leibe, and Ramanan], it often suffers from under-reconstruction [Kerbl et al.(2023)Kerbl, Kopanas, Leimkühler, and Drettakis] during the process of Gaussian densification [Zhang et al.(2024)Zhang, Zhan, Xu, Lu, and Xing], which affects both static and dynamic scenes. The under-reconstruction can be clearly observed with blur and artifacts in the rendered 2D images, the discrepancy of frequency spectrum of the render images and the corresponding ground truth, and the predicted optical flow results by [Xu et al.(2023)Xu, Zhang, Cai, Rezatofighi, Yu, Tao, and Geiger], as illustrated in Fig. 1.